# Assessing the Impact of Upselling in Online Fantasy Sports

### v1

@@ 1. Introduction
Upselling is studied extensively across multiple sectors, such as airlines, hospitality, retail, and insurance. And their impacts on customer satisfaction and revenue metrics. For instance, in the airline industry, upselling involves flight seat upgrades (Thirumuruganathan et al., 2023), while in the hospitality industry, it involves luxury room upgrades (Guillet, 2020),(Guelman et al., 2014) Insurance industry to cross-sell/Upsell.
In this research paper, we study user behaviour after the Dream11 business’s upselling and develop a system to optimise the upselling policy, which has not been effectively addressed by existing methods relying on traditional machine learning techniques.
@@ 3.1. Upselling Experiment
The payments dataset used in this research comes from millions of payment transactions conducted on the Dream11 fantasy sports gaming platform. Whenever the user adds cash to the account, we recommend a set of three values shown in table 2. There are maximum limits and UX constraints on these values. We arrived at these values by analyzing historical deposit amounts and identifying the most frequent deposit amounts and their multiples. This allowed us to cover a range of common deposit behaviours and encourage higher deposits without overwhelming the user. Below are
@@
The objective was to experiment with upselling by recommending higher deposit amounts with guardrails on the following user engagement metrics. Deposit funnel: Total users attempted to make a transaction -¿ Total users successfully made a transaction. User Conversion: Paid registrations /Total registrations Using an intensity factor ($a$) and the user deposit propensity values, we implemented the following grid experimentation strategy:
  1. (1)
    Sample deposit amounts from $\text{F}(X_{i})$ for each user.
  2. (2)
    Multiply intensity ($a$) to the prediction and arrive at the recommended values.
For each user ($i$), we randomly assigned the intensity treatments to 4 different target groups (TGs), as given in Table [3].
@@ 3.2. Online Results
The control group received a default recommendation, whereas the target groups received various personalized amounts alongside different discrete intensity values [1x, 1.25x, 1.5x, 2x]. This experiment lasted eight weeks, from January 1, 2024, to February 28, 2024. We observed an immediate impact on the per transaction value of about +5.6%, while users were also making fewer transactions on the application, resulting in a 4.8% fall in the most aggressive variant. While evaluating the long-term sustained impact in the avg_deposit_week_8, Figure [1] illustrates the uplift across the experimental variants shown in Table [3].
@@ 4. Causal Uplift Modelling
While analyzing the experiment of varying the intensity parameter, The challenge was to estimate how much to upsell a user without a trade-off in user satisfaction. A popular approach for efficiently assigning treatments to a subset of customers for targeted personalization is using causal machine learning models (Gutierrez and Gérardy, 2017; Diemert, 2018). For any causal uplift modelling task, we need to gather data on three essential items: Treatment (T), Response (Y), and Feature / Covariate (X).
@@ 4.1. Dataset Preparation
We split the data into two parts: data_train and data_test collected from the randomized control trials in Section 3
  • Treatment (Ti): Intensity [1.0, 1.25, 1.50, 2.0]—
  • Response Metric (Y): Deposit
  • Xi: User Features, Gameplay Features, User Demographics
We can arrive at an optimal upselling policy by personalising the intensity factor for each user, We estimate the incremental uplift after the intensity treatments (Ti) is given to each user uniformly. Next, we will use this dataset to model upselling.
@@ 4.2. Model Training
We estimated the CATE $Y(x)$, the conditional increment in the probability of completing the deposit $\Pr(Y=1)$ due to upselling intensity, given user features (X). We tested the following algorithms: S-learner, T-learner, X-learner, and R-learner, implemented in Python using (Chen et al., 2020), with XGBRegressor as the base learner and a propensity score of 0.5.
For training, we used the hyperopt package (Bergstra et al., 2013) with SparkTrials for distributing our workloads and managed multiple model runs with the MLflow library (Zaharia et al., 2018). We used the Tree-based Parzen Estimators (TPE) for efficient parameter search and defined the objective function as follows:
  • Initiate Models: Obtain models from causalml (Chen et al., 2020) and define the parameter search space.
  • Train Models on Training Data: Train the models using the training dataset.
  • Infer Models: Apply the trained models to the training and test datasets.
  • Create Validation/Optimization Metric: Calculate the mean of the test set AUUC (Area Under the Uplift Curve) scores for validation and optimization.
We conducted 1000 random trials, logging the AUUC for each treatment on the test set. We evaluated two model configurations: The global configuration, selecting the model with the lowest test set loss, and the local configuration, predicting the Conditional Average Treatment Effect (CATE) for the i-th treatment using the model with the highest test set i-th AUUC.
Table [4] summarises the configurations for the XGBRegressor used in the spark trials. Each treatment group (TG1, TG2, TG3, TG4) and the global model (TG_global) have tailored local configuration parameters, while the global configuration is optimized for overall performance, and local configurations are optimised for individual treatments. We compare the CATE scores for each treatment and keep a threshold of greater than zero else a control treatment is assigned to the user as shown below.
$\pi(x)=\begin{cases}\operatorname{argmax}\limits_{t\neq\text{CG}}\text{CATE}(x\mid T=t)&\text{if}\;\max\limits_{t\neq\text{CG}}\text{CATE}(x\mid T=t)>\text{0}\\ \text{CG}&\text{otherwise}\end{cases}$
@@
Table 4. XGBRegressor Configurations
@@ 4.3. Evaluation
    Treatment Grid exp policy CATE policy
@@
    We calculate the expected bootstrapped revenue with 5000 as the sample size. The CATE policies Test(pi) and Train(pi) arrive after model inference, and the intensity factor with the highest uplift is selected. For the CATE policy, Figure [2] shows a significant 10.7% uplift in both train and test datasets.
@@
  • AUUC There is an Uplift Curve and a random curve based on the calculation of every treatment. The Uplift charts indicate that there is good offline fitment of the underlying estimators, Cum gain above is plotted by ordering the users based on the inferred uplift scores, and we observe that most of the gains are captured in the initial set of users.
@@ 5.1. Conclusion
Our research demonstrates the complex trade-offs between upselling intensity and user satisfaction. We observed that increasing intensity in upselling strategies led to higher immediate revenue but decreased user recall and satisfaction. This was particularly evident in the TG1 group, where the bi-modal distribution of user responses indicates a polarized reaction to upselling efforts. These findings underscore the importance of balancing upselling intensity with user experience to ensure user engagement and revenue uplift.
The causal uplift modelling framework helps us determine a personalised intensity value for each user, balancing the trade-offs and ensuring fairness in the recommendations. The causal analysis results across both test and train datasets indicate that taking the models online can be promising, with an expected revenue uplift of 10.7%.
We further analysed the online results in Figure [1], the guardrails were majorly affected for new users in the system. The offline CATE policy treats with 1x intensity to 80% of new users, leading to an impressive 2.1% improvement in conversion rates. While the offline results are promising, we must still validate the optimal trade-offs in an online experimental setting.

### v2

@@ 1. Introduction
Upselling is studied extensively across multiple sectors, such as airlines, hospitality, retail, and insurance. And their impacts on customer satisfaction and revenue metrics. For instance, in the airline industry, upselling involves flight seat upgrades (Thirumuruganathan et al., 2023), while in the hospitality industry, it involves luxury room upgrades (Guillet, 2020), In the Insurance industry (Guelman et al., 2014) to cross-sell/Upsell.
Upselling applications span sectors. In this paper, we study user behaviour of upselling on a fantasy sports platform, Dream11, with over 200M users. Further, to optimise the upselling policy, we developed a system that has not been effectively addressed by existing methods relying on traditional machine learning techniques.
@@ 3.1. Upselling Experiment
The payments dataset used in this research comes from millions of payment transactions conducted on the Dream11 fantasy sports gaming platform. Whenever the user adds cash to the account, we recommend a set of three values shown in table 2. There are maximum limits and UX constraints on these values. We arrived at these values by analyzing historical deposit amounts and identifying the most frequent deposit amounts and their multiples. This allowed us to cover a range of common deposit behaviours and encourage higher deposits without overwhelming the user.
@@
3.1.1. Objective:
The aim was to evaluate the impact of upselling using an intensity factor ($a$). We sampled deposit amounts from the user-specific distribution $\text{F}(X_{i})$ and adjusted them using $a$. Users were randomly assigned to one of the four target groups (TGs) listed in Table 3.
@@ 3.1.1. Objective:
3.1.2. Guardrails
We monitored the following metrics to ensure upselling didn’t negatively affect user experience:
  • Conversion Funnel: Ratio of successful transactions to attempted transactions.
  • User Engagement: Recall on the recommendations, Time spent on the platform and changes in transaction frequency.
  • Revenue Uplift: Comparison of average deposit amounts across target groups.
@@ 3.2. Online Results
The control group received a default recommendation, whereas the target groups received various personalized amounts alongside different discrete intensity values [1x, 1.25x, 1.5x, 2x]. This experiment lasted eight weeks, from January 1, 2024, to February 28, 2024. We observed an immediate impact on the per transaction value of about +5.6%, while users were also making fewer transactions on the platform (reduction in infra cost), resulting in a 4.8% fall in the most aggressive variant. While evaluating the long-term sustained impact in the avg_deposit_week_8, Figure [1] illustrates the uplift across the experimental variants shown in Table [3].
@@ 4. Causal Uplift Modelling
During data analysis of the upselling experiment, we saw that the challenge was to estimate how much to upsell a user without a trade-off in user satisfaction. A popular approach for efficiently assigning treatments to a subset of customers for targeted personalization is using causal machine learning models (Gutierrez and Gérardy, 2017; Diemert, 2018).
@@ 4.1. Dataset Preparation
For causal uplift modeling, the dataset consists of three key components: Treatment (T), Response (Y), and Features/Covariates (X). These are defined as follows:
  • Treatment (Ti): The intensity of upselling, represented by four levels: {1.0, 1.25, 1.50, 2.0}.
  • Response (Y): The primary outcome of interest, measured as the total deposit amount by each user.
  • Features (Xi): User-specific characteristics, including demographics, gameplay behavior, and historical transaction data.
To arrive at an optimal upselling policy, we personalize the intensity factor for each user. We estimate the incremental uplift by applying the intensity treatments (Ti) uniformly across all users. This dataset will then be used to model and optimize upselling strategies.
@@ 4.2. Model Training
We estimated the Conditional Average Treatment Effect (CATE) $Y(x)$, which represents the change in the probability of completing a deposit $\Pr(Y=1)$ due to varying upselling intensity, based on user features $X$. To achieve this, we tested four algorithms: S-learner, T-learner, X-learner, and R-learner, implemented using Python’s causalml library (Chen et al., 2020), with XGBRegressor as the base model and a fixed propensity score of 0.5.
For model training and optimization, we used the hyperopt package (Bergstra et al., 2013) with SparkTrials to distribute workloads, and MLflow (Zaharia et al., 2018) to manage model runs. The parameter search was guided by Tree-based Parzen Estimators (TPE), with the objective function defined as:
  • Define the parameter search space using causalml.
  • Train models on the training dataset.
  • Apply trained models to both training and test sets.
  • Optimize by calculating the mean AUUC (Area Under the Uplift Curve) score on the test set.
We ran 1000 random trials, logging the AUUC for each treatment. Two model configurations were tested:
  • Global configuration: Selects the model with the lowest test set loss across all treatments.
  • Local configuration: Optimizes the CATE for each treatment group (TG) using the highest test set AUUC.
Table 4 summarizes the XGBRegressor configurations used in the SparkTrials. Each treatment group (TG1, TG2, TG3, TG4) has its own local configuration parameters, while the global model (TG_global) is optimized for overall performance. A treatment is assigned based on the following rule:
$\pi(x)=\begin{cases}\operatorname{argmax}\limits_{t\neq\text{CG}}\text{CATE}(x\mid T=t)&\text{if}\;\max\limits_{t\neq\text{CG}}\text{CATE}(x\mid T=t)>0\\ \text{CG}&\text{otherwise}\end{cases}$
@@
Table 4. XGBRegressor Configurations for Each Treatment Group
@@ 4.3. Evaluation
    Treatment Grid exp policy CATE policy($\pi$)
@@
    We calculate the expected bootstrapped revenue using a sample size of 5000. After model inference, the CATE policies for both test ($\pi_{\text{test}}$) and train ($\pi_{\text{train}}$) datasets are derived, with the intensity factor yielding the highest uplift being selected. As shown in Figure [2], the CATE policy results in a significant 10.7% uplift across both the training and test datasets.
@@
  • AUUC (Area Under the Uplift Curve): The uplift chart consists of an Uplift Curve and a random baseline curve for each treatment. The charts demonstrate a strong offline fit of the underlying estimators. The cumulative gains are plotted by ranking users based on their inferred uplift scores, showing that most gains are concentrated within the top-ranked users.
@@ 5.1. Conclusion
Our research highlights the intricate trade-offs between upselling intensity and user satisfaction. While increasing upselling intensity leads to immediate revenue gains, it also results in a decline in user recall and satisfaction. This effect was most pronounced in the TG1 group, where a bi-modal distribution of responses revealed polarized reactions to upselling strategies. These findings emphasize the need to carefully balance upselling intensity with user experience to maintain both engagement and revenue growth.
The causal uplift modelling framework allows us to tailor the intensity for each user, effectively managing these trade-offs and ensuring fairness in recommendations. Results from both the test and training datasets show that taking these models online holds significant potential, with an estimated revenue uplift of 10.7%.
Further analysis, as shown in Figure [1], revealed that the guardrails mainly were impacted for new users. The offline CATE policy applied a 1x intensity for 80% of new users, resulting in a 2.1% improvement in conversion rates. While these offline results are promising, further validation is required in an online experimental setting to confirm the optimal trade-offs between revenue and user experience.